you are creating a rectangular banner for a school pep rally. you have 100 feet squared of paper and you want the length to be 15 ft longer than the width. what should be the demensions of the banner be?
Answer by oberobic(2304)

Area = Length * Width
A = L * W
A = 100 sq. ft
W = ?
L = W + 15
Substituting
100 = (W+15)*W
100 = W^2+15W
Subtract 100 from both sides
0 = W^2 + 15W -100
Factor
What are the relevant factors of 100, one positive and one negative that when added equal 15?
5 * 20 = 100
20 -5 = 15
OK
(W-5)(W+20) = 0
So,
W = 5
W = -20
But
We cannot have negative width, so our only applicable solution is
W = 5
Looking back, we see that
L = W + 15
L = 5 + 15
L = 20
Answer:
The banner will have length of 20 and width of 5.
Done.